# Optomechanical self-organization in a mesoscopic atom array

### v1

@@ II Mesoscopic signatures of a phase transition
To probe the evolution of system properties with size directly, we examine how critical behavior varies with atom number. We first examine the dependence of $\mathrm{\Omega_{c}}$ on $N$, and find that $\Omega_{c}$ diminishes as $N$ increases. Quantitatively, for an averaging time of $5\,\mu\mathrm{s}$, we almost recover the $\sim 1/\sqrt{N}$ dependence predicted by Equation (4), shown in Fig. 2B. Though the critical point’s scaling with $N$ has been measured before in macroscopic bosonic and fermionic realizations of the superradiant phase transition [43, 47], our work shows that the $N$-dependence is also observable in a regime far from the thermodynamic limit. In Fig. 2B, the data appear above the prediction of Equation (4) if we use the temperature from our independent measurement, implying that the atoms have been heated to a temperature close to $66\,\mu$K at the time of observation.
@@
Figure 3: Lifetime of the self-organized state. Above the critical point, we observe three dynamical behaviors in the cavity field time traces: (A) symmetry breaking, (B) switching, and (C) rapid oscillations. Time is measured from the end of the pump ramp-up. Shaded areas show the time spent in the $c_{\mathrm{proj}}>0$ state (red) and the $c_{\mathrm{proj}}<0$ state (blue). Shaded purple regions indicate the shot noise level. Data are shown for $N=20,\,\mathrm{\Omega}=2\pi\times 45.4\,\mathrm{MHz}$. (D) $g^{(1)}(0,t)$ for $N=20$ at $\mathrm{\Omega}<\mathrm{\Omega_{c}}$ and $\mathrm{\Omega}>\mathrm{\Omega_{c}}$. When the system is self-organized, $g^{(1)}(0,t)$ decays on a much longer timescale. All data in figure are taken at $\mathrm{\Delta_{pa}}=-2\pi\times 80\,\mathrm{MHz}$ and $\mathrm{\Delta_{pc}}=-2\pi\times 1.9\,\mathrm{MHz}$. (E) Dependence of the critical point on $N$ for different averaging times. At larger $N$, the curves converge, indicating that self-organization persists on longer timescales. At lower $N$, the separation between the curves grows due to the transience of the mesoscopic system. The measurements at $25\,\mu$s and $50\,\mu$s for $N<14$ are absent because the transition was not reached in the range of $\mathrm{\Omega}$ values scanned. The plotted $\Omega_{\mathrm{c}}$ data are subject to the same systematic errors as described in Fig. 2B. Lines are guides to the eye.
@@
Atomic self-organization in a cavity can be compared with the crossover between a magnetized and unmagnetized state in a spin system. Notably, due to the infinite-range character of the interaction, the crossover does not occur by forming domains of increasing size as the critical coupling $\Omega_{\mathrm{c}}$ is approached [42]. Instead, the behavior of $\langle\hat{\Theta}\rangle$ above the transition is analogous to that of a superparamagnet: As discussed in seminal work by Néel [48], small ferromagnetic particles are subject to thermal fluctuations that cause their magnetization to relax on a time scale that grows with the particle volume. Averaging the magnetization over times much longer than the relaxation time results in zero net magnetization, while averaging over times much shorter than the relaxation time allows for observation of quasistatic magnetic ordering [5]. We similarly find that mesoscopic self-organization of atoms has a characteristic lifetime that depends on system size.
@@
To reveal the dynamics of the magnetization order parameter $\langle\hat{\mathrm{\Theta}}\rangle$, we take advantage of the cavity’s $\mu$s-scale readout capability [35]. In Fig. 3, we examine single-shot time traces of $c_{\mathrm{proj}}$, time correlations of the field, as well as the $N$-dependence of the critical point as we increase the averaging time. Fig. 3A-C show single-shot time traces representative of observed dynamical behaviors in the self-organized regime: symmetry breaking, switching, and rapid oscillations. For a macroscopic atom number, only symmetry breaking should occur, while for a mesoscopic atom number, fluctuations away from equilibrium can lead to switching between stable states of the system [42]. The rapid oscillation events are rare, occurring in only 2 out of the 180 shots obtained at the shown parameter setting; we expect they result from the energy in the dominant mode being above the energy gap separating the two symmetry-broken states.
@@
To highlight the difference in dynamical behavior below and above the critical point, we calculate the first-order coherence function of the field quadrature $g^{(1)}(t_{1},t_{2})$, defined as $g^{(1)}(t_{1},t_{2})=\langle c^{*}_{\mathrm{proj}}(t_{1})c_{\mathrm{proj}}(t_{2})\rangle/\sqrt{\langle|c_{\mathrm{proj}}(t_{1})|^{2}\rangle\langle|c_{\mathrm{proj}}(t_{2})|^{2}\rangle}$, where the angle brackets denote averaging over multiple shots. In particular, $g^{(1)}(0,t)$ indicates how correlated the field at time $t_{2}=t$ is to the field at time $t_{1}=0$, where $t_{1}=0$ is chosen to be at the end of the pump ramp-up. Symmetry breaking is indicated by $g^{(1)}(0,t)$ remaining high for extended $t$. In Fig. 3D, we see that when $\mathrm{\Omega}/\mathrm{\Omega_{c}}=0.8$, $g^{(1)}(0,t)$ decays rapidly, indicating the absence of symmetry breaking, and exhibits oscillations at approximately $60\,\mathrm{kHz}$, corresponding to the softened trap frequency. At $\mathrm{\Omega}/\mathrm{\Omega_{c}}=1.6$, $g^{(1)}(0,t)$ decays on a much longer timescale and oscillations are absent, indicating a prolonged lifetime of the symmetry-broken states.
@@
Figure 4: Optomechanical susceptibility. (A) Schematic of the susceptibility measurement. The dominant mode of the array ($z_{\mathrm{dom}}$) is displaced by $\delta z$. When $\delta z=0$, the effective potential $V_{\mathrm{eff}}$ on the atoms is symmetric about $z_{\mathrm{dom}}=0$ (i). For $|\delta z|>0$, $V_{\mathrm{eff}}$ tilts in the direction of the bias (ii-iii). (B) Distributions of $c_{\mathrm{proj}}/\bar{c}_{\mathrm{a.n.}}$ (small purple points) and their mean values (circles) at various $\delta z$ for $N=18,\,\mathrm{\Omega/\Omega_{c}}=1.16$. Measurement of $\partial(\bar{c}_{\mathrm{proj}}/\bar{c}_{\mathrm{a.n.}})/\partial(\delta z)\big{\rvert}_{\delta z=0}$ is taken by fitting the slope of $\bar{c}_{\mathrm{proj}}/\bar{c}_{\mathrm{a.n.}}$ for $-0.01\lambda\leq\delta z\leq 0.01\lambda$ (red circles). (C) Susceptibility $\chi$ at various $\mathrm{\Omega}/\mathrm{\Omega_{c}}$ for $N=14$ and $N=18$. The data (hexagons and circles) are plotted together with the theoretical prediction based on the Boltzmann distribution (dashed lines). All data in figure are taken at $\mathrm{\Delta_{pa}}=-2\pi\times 80\,\mathrm{MHz}$ and $\mathrm{\Delta_{pc}}=-2\pi\times 1.9\,\mathrm{MHz}$.
@@ III Finite susceptibility at criticality
Here, $\bar{c}_{\mathrm{a.n.}}$ is the measurement of $\bar{c}_{\mathrm{proj}}$ when the tweezers are positioned at the positive antinodes. The overbar indicates an average over the distribution of $c_{\mathrm{proj}}$ obtained from the first $50\,\mu\mathrm{s}$ of multiple shots. The quantity $\partial(\bar{c}_{\mathrm{proj}}/\bar{c}_{\mathrm{a.n.}})/\partial(\delta z)\big{\rvert}_{\delta z=0}$ is measured by fitting the slope of $\bar{c}_{\mathrm{proj}}/\bar{c}_{\mathrm{a.n.}}$ in the range $-0.01\lambda\leq\delta z\leq 0.01\lambda$ for $N=18$ (red circles in Fig. 4B) and $N=14$. We note that the atoms are initialized with root-mean-square thermal fluctuations in position of about 100 nm, leading to a thermal Lamb-Dicke parameter of about 0.8 [45].
@@
From measurement of $\chi$ for $N=14$ and $N=18$ at different values of $\mathrm{\Omega}/\mathrm{\Omega_{c}}$ (Fig. 4C), we make two observations: 1) while $\chi$ increases with $\mathrm{\Omega/\Omega_{c}}$, it does not diverge at the critical point as one would expect for a second-order phase transition in a macrosopic system, and 2) the value of $\chi$ grows with atom number for $\mathrm{\Omega/\Omega_{c}}>1$. These observations can be explained by a model assuming $\chi\simeq\partial\langle z_{\mathrm{dom}}\rangle/\partial(\delta z)\big{\rvert}_{\delta z=0}=NM\nu^{2}\langle\Delta z_{\mathrm{dom}}^{2}\rangle/(k_{B}T)\big{\rvert}_{\delta z=0}$, where angle brackets indicate a thermal average including the tweezer bias $\delta z$[45]. Thus, the susceptibility is directly related to the size of fluctuations in the system through the variance $\langle\mathrm{\Delta}z_{\mathrm{dom}}^{2}\rangle$. In a macroscopic system, one would expect $\chi$ to diverge according to $1/(1-(\mathrm{\Omega/\Omega_{c}})^{2})$ as $\mathrm{\Omega}$ approaches $\mathrm{\Omega_{c}}$. Instead, because of our finite atom number, our model predicts that $\chi$ should remain finite at the critical point.
@@
Specifically, at the critical point, our model predicts that $\chi$ should be approximately proportional to $\sqrt{N}$ as a result of the variance $\langle\mathrm{\Delta}z_{\mathrm{dom}}^{2}\rangle$ scaling as $1/\sqrt{N}$ at leading order. This has a simple interpretation: At smaller $N$, the fluctuations in $z_{\mathrm{dom}}$ are larger and the atoms sample the quartic part of the effective optomechanical potential $V_{\mathrm{eff}}$, which provides the confinement that prevents the susceptibility from diverging. While we observe finite $\chi$ at the critical point, we are unable to resolve the $\sqrt{N}$-dependence due to the quality of our data. Nonetheless, we emphasize that this observation of finite $\chi$ is a direct measure of mesoscopic fluctuations and not just a result of non-zero temperature—a macroscopic system at non-zero temperature would still exhibit a diverging susceptibility. As $\mathrm{\Omega}/\mathrm{\Omega_{c}}$ is increased beyond 1, our model predicts that $\chi$ should continue to grow and eventually scale proportionally to $N$, before saturating at a value of $NM\nu^{2}\lambda^{2}/(16k_{B}T)$. For $\mathrm{\Omega}\gg\mathrm{\Omega_{c}}$, the $\delta z$ required to confine $\langle z_{\mathrm{dom}}\rangle$ effectively to one of the minima of $V_{\mathrm{eff}}$ scales inversely with $N$, leading to $\chi\propto N$. We note that $\chi$ continuing to increase for $\Omega>\Omega_{\mathrm{c}}$ is a consequence of us applying the tweezer shift $\delta z$ to break symmetry before the cavity interaction is turned on, analogous to applying a magnetic field to a spin system before lowering its temperature below the Curie point. In contrast, if the tweezer shift were applied after the cavity interaction, one might expect $\chi$ to peak at $\Omega=\Omega_{\mathrm{c}}$. Accordingly, we see the separation between the $N=14$ and $N=18$ data become more pronounced as we increase $\mathrm{\Omega}$ beyond $\mathrm{\Omega_{c}}$, consistent with the prediction of $\chi$ approaching linear scaling with $N$. These results highlight how the response of mesoscopic systems deviates from that of macroscopic systems.
@@ IV Discussion and outlook
We have studied self-organization, bifurcation, and symmetry breaking as well as mesoscopic dynamics and response due to light-mediated interactions in an atom array. The deterministic atom number and precise control over interactions necessary for this work were achieved by interfacing an optical tweezer array with an optical cavity, opening new directions for these systems. The realization of mesoscopic self-organization with control over individual atoms has significant implications not just for our fundamental understanding of mesoscopic systems, but also for quantum simulation, optimization, and metrology. On one hand, the dynamics of the cavity field provide a plethora of information that may further our understanding of nonequilibrium physics in mesoscopic systems [49] and pave the way for time-resolved quantum simulation [50, 51]. On the other hand, mesoscopic self-organization may also open the door to programmable engineering [52] of exotic symmetry-breaking phases of matter [53], for instance by illuminating groups of atoms with different phases of pump light. This system can also be used to implement a quantum numerical solver for a variety of classically hard computational problems [54, 55, 56], where the tweezer array allows for highly programmable initial conditions. Additionally, the optomechanical susceptibility near the critical point can be a resource for enhanced sensing of applied forces [57, 58, 59].

### v2

@@ II Mesoscopic signatures of a phase transition
We now examine how critical behavior varies with atom number. We first examine the dependence of $\mathrm{\Omega_{c}}$ on $N$, and find that $\Omega_{c}$ diminishes as $N$ increases. Quantitatively, for an averaging time of $5\,\mu\mathrm{s}$, we almost recover the $\sim 1/\sqrt{N}$ dependence predicted by Equation (4), shown in Fig. 2b. Though the critical point’s scaling with $N$ has been measured before in macroscopic bosonic and fermionic realizations of the superradiant phase transition [41, 44], our work shows that the $N$-dependence is also observable in a regime far from the thermodynamic limit. In Fig. 2b, the data appear above the prediction of Equation (4) if we use the temperature from our independent measurement, implying that the atoms have been heated to a temperature close to $66\,\mu$K at the time of observation.
@@
Figure 3: Lifetime of the self-organized state. Above the critical point, we observe three dynamical behaviors in the cavity field time traces: a, symmetry breaking, b, switching, and c, rapid oscillations. Time is measured from the end of the pump ramp-up. Shaded areas show the time spent in the $c_{\mathrm{proj}}>0$ state (red) and the $c_{\mathrm{proj}}<0$ state (blue). Shaded purple regions indicate the shot noise level. Data are shown for $N=20,\,\mathrm{\Omega}=2\pi\times 45.4\,\mathrm{MHz}$. d, $g^{(1)}(0,t)$ for $N=20$ at $\mathrm{\Omega}<\mathrm{\Omega_{c}}$ and $\mathrm{\Omega}>\mathrm{\Omega_{c}}$. When the system is self-organized, $g^{(1)}(0,t)$ decays on a much longer timescale. e, Dependence of the critical point on $N$ for different averaging times. At larger $N$, the curves converge, indicating that self-organization persists on longer timescales. At lower $N$, the separation between the curves grows due to the transience of the mesoscopic system. The measurements at $25\,\mu$s and $50\,\mu$s for $N<14$ are absent because the transition was not reached in the range of $\mathrm{\Omega}$ values scanned. The plotted $\Omega_{\mathrm{c}}$ data are subject to the same systematic errors as described in Fig. 2b. Lines are guides to the eye. All data in figure are taken at $\mathrm{\Delta_{pa}}=-2\pi\times 80\,\mathrm{MHz}$ and $\mathrm{\Delta_{pc}}=-2\pi\times 1.9\,\mathrm{MHz}$.
@@
Atomic self-organization in a cavity can be compared with the crossover between a magnetized and unmagnetized state in a spin system [40]. The behavior of $\langle\hat{\Theta}\rangle$ above the transition is analogous to that of a superparamagnet: As discussed in seminal work by Néel [45], small ferromagnetic particles are subject to thermal fluctuations that cause their magnetization to relax on a time scale that grows with the particle volume. Averaging the magnetization over times much longer than the relaxation time results in zero net magnetization, while averaging over times much shorter than the relaxation time allows for observation of quasistatic magnetic ordering [46]. We similarly find that mesoscopic self-organization of atoms has a characteristic lifetime that depends on system size.
@@
To reveal the dynamics of the magnetization order parameter $\langle\hat{\mathrm{\Theta}}\rangle$, we take advantage of the cavity’s $\mu$s-scale readout capability [33]. Fig. 3a-c show single-shot time traces representative of observed dynamical behaviors in the self-organized regime: symmetry breaking, switching, and rapid oscillations. For a macroscopic atom number, only symmetry breaking should occur, while for a mesoscopic atom number, fluctuations away from equilibrium can lead to switching between stable states of the system [40]. The rapid oscillation events are rare, occurring in only 2 out of the 180 shots obtained at the shown parameter setting; we expect they result from the energy in the dominant mode being above the energy gap separating the two symmetry-broken states.
@@
To highlight the difference in dynamical behavior below and above the critical point, we calculate the first-order coherence function of the field quadrature $g^{(1)}(t_{1},t_{2})$, defined as $g^{(1)}(t_{1},t_{2})=\langle c^{*}_{\mathrm{proj}}(t_{1})c_{\mathrm{proj}}(t_{2})\rangle/\sqrt{\langle|c_{\mathrm{proj}}(t_{1})|^{2}\rangle\langle|c_{\mathrm{proj}}(t_{2})|^{2}\rangle}$, where the angle brackets denote averaging over multiple shots. In particular, $g^{(1)}(0,t)$ indicates how correlated the field at time $t_{2}=t$ is to the field at time $t_{1}=0$, where $t_{1}=0$ is chosen to be at the end of the pump ramp-up. Symmetry breaking is indicated by $g^{(1)}(0,t)$ remaining high for extended $t$. In Fig. 3d, we see that when $\mathrm{\Omega}/\mathrm{\Omega_{c}}=0.8$, $g^{(1)}(0,t)$ decays rapidly, indicating the absence of symmetry breaking, and exhibits oscillations at approximately $60\,\mathrm{kHz}$, corresponding to the softened trap frequency. At $\mathrm{\Omega}/\mathrm{\Omega_{c}}=1.6$, $g^{(1)}(0,t)$ decays on a much longer timescale and oscillations are absent, indicating a prolonged lifetime of the symmetry-broken states. While we are able to observe these qualitative differences in the decay of $g^{(1)}(0,t)$ for $\Omega<\Omega_{\mathrm{c}}$ and $\Omega>\Omega_{\mathrm{c}}$, we note that using $g^{(1)}(0,t)$ to determine the location of $\Omega_{\mathrm{c}}$ does not appear to be a robust method.
@@
Figure 4: Optomechanical susceptibility. a, Schematic of the susceptibility measurement. The dominant mode of the array ($z_{\mathrm{dom}}$) is displaced by $\delta z$. When $\delta z=0$, the effective potential $V_{\mathrm{eff}}$ on the atoms is symmetric about $z_{\mathrm{dom}}=0$ (i). For $|\delta z|>0$, $V_{\mathrm{eff}}$ tilts in the direction of the bias (ii-iii). b, Distributions of $c_{\mathrm{proj}}/\bar{c}_{\mathrm{a.n.}}$ (small purple points) and their mean values (circles) at various $\delta z$ for $N=18,\,\mathrm{\Omega/\Omega_{c}}=1.16$. Measurement of $\partial(\bar{c}_{\mathrm{proj}}/\bar{c}_{\mathrm{a.n.}})/\partial(\delta z)\big{\rvert}_{\delta z=0}$ is taken by fitting the slope of $\bar{c}_{\mathrm{proj}}/\bar{c}_{\mathrm{a.n.}}$ for $-0.01\,\lambda\leq\delta z\leq 0.01\,\lambda$ (red circles). c, Susceptibility $\chi$ at various $\mathrm{\Omega}/\mathrm{\Omega_{c}}$ for $N=14$ and $N=18$. The data (hexagons and circles) are plotted together with the theoretical prediction based on the Boltzmann distribution for $T=35\,\mu$K (dashed lines). All data in figure are taken at $\mathrm{\Delta_{pa}}=-2\pi\times 80\,\mathrm{MHz}$ and $\mathrm{\Delta_{pc}}=-2\pi\times 1.9\,\mathrm{MHz}$.
@@ III Finite susceptibility at criticality
Here, $\bar{c}_{\mathrm{a.n.}}$ is the measurement of $\bar{c}_{\mathrm{proj}}$ when the tweezers are positioned at the positive antinodes. The overbar indicates an average over the distribution of $c_{\mathrm{proj}}$ obtained from the first $50\,\mu\mathrm{s}$ of multiple shots. The quantity $\partial(\bar{c}_{\mathrm{proj}}/\bar{c}_{\mathrm{a.n.}})/\partial(\delta z)\big{\rvert}_{\delta z=0}$ is measured by fitting the slope of $\bar{c}_{\mathrm{proj}}/\bar{c}_{\mathrm{a.n.}}$ in the range $-0.01\,\lambda\leq\delta z\leq 0.01\,\lambda$ for $N=18$ (red circles in Fig. 4b) and $N=14$. We note that the atoms are initialized with root-mean-square thermal fluctuations in position of about 100 nm, which is comparable to $k^{-1}$, leading to a thermal Lamb-Dicke parameter of about 0.8 (see Supplementary Equation (S18)).
@@
From measurement of $\chi$ for $N=14$ and $N=18$ at different values of $\mathrm{\Omega}/\mathrm{\Omega_{c}}$ (Fig. 4c), we make two observations: 1) while $\chi$ increases with $\mathrm{\Omega/\Omega_{c}}$, it does not diverge at the critical point as one would expect for a second-order phase transition in a macrosopic system, and 2) the value of $\chi$ grows with atom number for $\mathrm{\Omega/\Omega_{c}}>1$. These observations can be explained by a model assuming $\chi\simeq\partial\langle z_{\mathrm{dom}}\rangle/\partial(\delta z)\big{\rvert}_{\delta z=0}=NM\nu^{2}\langle\Delta z_{\mathrm{dom}}^{2}\rangle/(k_{B}T)\big{\rvert}_{\delta z=0}$, where angle brackets indicate a thermal average including the tweezer bias $\delta z$ (see Supplementary Information Section IV). Thus, the susceptibility is directly related to the size of fluctuations in the system through the variance $\langle\mathrm{\Delta}z_{\mathrm{dom}}^{2}\rangle$. In a macroscopic system, one would expect $\chi$ to diverge according to $1/(1-(\mathrm{\Omega/\Omega_{c}})^{2})$ as $\mathrm{\Omega}$ approaches $\mathrm{\Omega_{c}}$. Instead, because of our finite atom number, our model predicts that $\chi$ should remain finite at the critical point.
@@
Specifically, at the critical point, our model predicts that $\chi$ should be approximately proportional to $\sqrt{N}$ as a result of the variance $\langle\mathrm{\Delta}z_{\mathrm{dom}}^{2}\rangle$ scaling as $1/\sqrt{N}$ at leading order. This has a simple interpretation: At smaller $N$, the fluctuations in $z_{\mathrm{dom}}$ are larger and the atoms sample the quartic part of the effective optomechanical potential $V_{\mathrm{eff}}$, which provides the confinement that prevents the susceptibility from diverging. While we observe finite $\chi$ at the critical point, we are unable to resolve the $\sqrt{N}$-dependence due to the quality of our data. Nonetheless, we emphasize that this observation of finite $\chi$ is a direct measure of mesoscopic fluctuations and not just a result of non-zero temperature—a macroscopic system at non-zero temperature would still exhibit a diverging susceptibility. As $\mathrm{\Omega}/\mathrm{\Omega_{c}}$ is increased beyond 1, our model predicts that $\chi$ should continue to grow and eventually scale proportionally to $N$, before saturating at a value of $NM\nu^{2}\lambda^{2}/(16k_{B}T)$. For $\mathrm{\Omega}\gg\mathrm{\Omega_{c}}$, the $\delta z$ required to confine $\langle z_{\mathrm{dom}}\rangle$ effectively to one of the minima of $V_{\mathrm{eff}}$ scales inversely with $N$, leading to $\chi\propto N$. Our model assumes that the atom temperature is the same for all $N$; while recoil heating may contribute to an $N$-dependent temperature and in turn affect the scaling of $\chi$ with $N$, the $1/\sqrt{N}$ scaling of $\Omega_{\mathrm{c}}$ observed in Fig. 2b belies any strong $N$-dependence of the temperature. We note that $\chi$ continuing to increase for $\Omega>\Omega_{\mathrm{c}}$ is a consequence of us applying the tweezer shift $\delta z$ to break symmetry before the cavity interaction is turned on, analogous to applying a magnetic field to a spin system before lowering its temperature below the Curie point. In contrast, if the tweezer shift were applied after the cavity interaction, one might expect $\chi$ to peak at $\Omega=\Omega_{\mathrm{c}}$. Accordingly, we see the separation between the $N=14$ and $N=18$ data become more pronounced as we increase $\mathrm{\Omega}$ beyond $\mathrm{\Omega_{c}}$, consistent with the prediction of $\chi$ approaching linear scaling with $N$. These results highlight how the response of mesoscopic systems deviates from that of macroscopic systems. Furthermore, the use of tweezers to displace the atoms provides a novel method for measuring optomechanical susceptibility, complementing experiments where the external perturbation is introduced by injecting a weak field into the cavity [47, 48].
@@ IV Discussion and outlook
We have studied self-organization, bifurcation, and symmetry breaking as well as mesoscopic dynamics and response due to cavity-mediated interactions in an atom array, opening new directions for these systems. The realization of mesoscopic self-organization with control over individual atoms has significant implications not just for our fundamental understanding of mesoscopic systems, but also for quantum simulation, optimization, and metrology. On one hand, the dynamics of the cavity field provide a plethora of information that may further our understanding of finite-size scaling [49] and nonequilibrium physics in mesoscopic systems [50], as well as pave the way for time-resolved quantum simulation [51, 52]. On the other hand, mesoscopic self-organization may also open the door to programmable engineering [53] of exotic symmetry-breaking phases of matter [54], for instance by illuminating groups of atoms with different phases of pump light. This system can also be used to implement a quantum numerical solver for a variety of classically hard computational problems [55, 56, 57], where the tweezer array allows for highly programmable initial conditions. Additionally, the optomechanical susceptibility near the critical point can be a resource for enhanced sensing of applied forces [58, 59, 60].